OpenStudy (anonymous):

f(x) = -3x -7, g(x) = -8x^2 -7x +8 find (F o G)(-7)

OpenStudy (anonymous):

@zepdrix I should be able to get this one without much trouble haha

OpenStudy (anonymous):

I think I need to rewrite it as f(g(-7)) first

OpenStudy (anonymous):

and then is it -3(-7) -7 = 14

jimthompson5910 (jim_thompson5910):

g(-7) is what

OpenStudy (anonymous):

g(x), x=-7

jimthompson5910 (jim_thompson5910):

yes what is g(x) when x = -7

jimthompson5910 (jim_thompson5910):

g(x) = -8x^2 -7x +8 g(-7) = -8(-7)^2 -7(-7) +8 g(-7) = ???

OpenStudy (anonymous):

oh, I plug in the f(x)?

jimthompson5910 (jim_thompson5910):

what answer do you get for g(-7)

OpenStudy (anonymous):

-335

jimthompson5910 (jim_thompson5910):

that's the input for f(x) this is because (F o G)(-7) is really f(g(-7))

jimthompson5910 (jim_thompson5910):

f(g(-7)) = f(-335) = ???

OpenStudy (anonymous):

998!

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks you!

jimthompson5910 (jim_thompson5910):

yw